Ms. Ling has 14 boys and 17 girls in her third-grade
class. How many ways can she select 2 boys or 2 girls
to make a Thanksgiving poster?

Answer by Aldorozos(172)

You need to know about combination http://betterexplained.com/articles/easy-permutations-and-combinations/
Here we have the option of either boys or girls. Therefore we have to add both options together.
for boys we have 14!/((14-2)!*2! = 91
for girls we have 17!/(17-2)!* 2! = 136
Now we have to add our two options together = 91+136